Session: Jun 2011 (session 2)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Yebes (Paco Colomer at Wed Jun 01 20:22:21 2011) Success:
No known problems.

Shanghai (Bo Xia at Wed Jun 01 21:35:25 2011) Success:

Urumqi (Wenjun Yang at Thu Jun 02 02:24:06 2011) Success:

Westerbork (Geert Kuper at Fri Jun 03 06:58:22 2011) Success:
No known problems.

Torun (K. Borkowski at Fri Jun 03 07:37:33 2011) Minor failures: WIND HIGH
Break in observations between 13:39 and 16:29 due to thunderstorm. One of claps of thunder was associated with a short loss of power which resulted in BBC disconnection. BBC04 could not be forced to the 'locked' state. With high Tsys probably whole this observation will be of little use for the PI.

Medicina (Andrea Orlati at Fri Jun 03 09:25:12 2011) Failure: RECO
Mark5 fails to record at 1Gbit/sec

Onsala (Michael Lindqvist at Mon Jun 06 05:15:10 2011) Success:
Tsys IF2 high. Otherwise no known problems.

Effelsberg (Uwe Bach at Mon Jun 06 09:47:50 2011) Minor failures: RECO
First scan lost because of Mark5Server crashed. Recording okay after 12:58 UT